# Early-Exit Graph Neural Networks

### v1

@@ 4.1 Impact of the Early-Exit Components
Model Amazon Ratings Minesweeper Roman Empire Tolokers Questions ACC $\uparrow$ AUROC $\uparrow$ ACC $\uparrow$ AUROC $\uparrow$ AUROC $\uparrow$ GCN 53.80 ± 0.60 97.86 ± 0.52 91.27 ± 0.20 83.64 ± 0.67 79.02 ± 1.27 SAGE 55.40 ± 0.21 97.77 ± 0.62 91.06 ± 0.67 82.43 ± 0.44 77.21 ± 1.28 GAT 55.54 ± 0.51 97.73 ± 0.73 90.63 ± 0.14 83.78 ± 0.43 77.95 ± 0.51 GT 51.17 ± 0.66 91.85 ± 0.76 86.51 ± 0.73 83.23 ± 0.64 77.95 ± 0.68 Polynormer 54.81 ± 0.49 97.46 ± 0.36 92.55 ± 0.37 85.91 ± 0.74 78.92 ± 0.89 Co-GNN 54.17 ± 0.37 97.31 ± 0.41 91.37 ± 0.35 84.45 ± 1.17 76.54 ± 0.95 SAS-GNNnoedge 51.47 ± 0.68 93.29 ± 0.61 83.46 ± 0.61 85.80 ± 0.79 79.60 ± 1.15 EEGNN 51.47 ± 0.51 93.18 ± 1.37 80.36 ± 0.43 85.26 ± 0.65 78.90 ± 1.15
Table 4: LRGB datasets. The scores are marked in red for the first, blue for the second, and green for the third.
Model Peptides-func Peptides-struct Pascal VOC-SP AP $\uparrow$ MAE $\downarrow$ F1 $\uparrow$ Classic MPNNs GCN 68.60±0.50 0.2460±0.0013 20.78±0.31 GINE 66.21±0.67 0.2473±0.0017 27.18±0.54 GatedGCN 67.65±0.47 0.2477±0.0009 38.80±0.40 GTs GT+LapPE 63.26±1.26 0.2529±0.0016 26.94±0.98 SAN+LapPE 63.84±1.21 0.2683±0.0043 32.30±0.39 GraphGPS+LapPE 65.35±0.41 0.2500±0.0005 37.48±1.09 Asynchronous MPNNs Co-GNNs 69.90 ± 0.93 - - AMP 71.63 ± 0.58 0.2431 ± 0.0004 - Ours SAS-GNNnoedge 69.71 ± 0.62 0.2449 ± 0.0013 22.65 ± 0.27 SAS-GNNedge 69.27 ± 0.58 0.2547 ± 0.0163 23.31 ± 0.49 SAS-GNNours 69.44 ± 0.63 0.2528 ± 0.0130 25.64 ± 0.63 EEGNNours 68.23 ± 0.37 0.2532 ± 0.0050 24.10 ± 0.73
4.2 Comparison with the Heterophilic and LRGB Benchmarks
We evaluate EEGNN on two widely studied benchmark settings. To isolate the impact of the early-exit mechanism, we also study SAS-GNN as an ablation of EEGNN with a fixed-depth architecture. For instance, SAS-GNN required tuning the number of layers $L$ for each dataset.
Datasets. We consider two evaluation settings. First, we use the heterophilic node classification benchmarks from Platonov et al. (2023). Second, we evaluate on three LRGB datasets: Peptides-func (graph classification), Peptides-struct (graph regression), and Pascal-VOC (inductive node classification). Additional results on transductive homophilic datasets and short-range graph classification tasks from the TUDataset collection (Morris et al., 2020) are presented in Appendix F.7 and F.8. Full dataset and implementation details are available in Appendix E.
Baselines. For the heterophilic benchmarks, we follow Platonov et al. (2023), comparing against GCN, GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GT (Dwivedi and Bresson, 2021a)—simple models that have proven effective even in heterophilic settings. We also include Co-GNN, a representative asynchronous MPNN, and Polynormer, a graph Transformer. For the LRGB benchmarks, we group baselines into: (i) classical MPNNs such as GCN, GINE (Hu et al., 2020), and GatedGCN (Bresson and Laurent, 2018); (ii) graph Transformers including GT, SAN (Kreuzer et al., 2021), and GraphGPS (Rampášek et al., 2023); and (iii) asynchronous MPNNs such as Co-GNN and AMP.
Results. EEGNN and SAS-GNN, implemented as in Algorithm 1 without edge features and under Corollary 3.3, perform competitively across both heterophilic datasets (Table 3) and LRGB benchmarks (Table 4). Despite their simplicity, lacking normalization or dropout, SAS-GNN ranks in the top two on 2 out of 5 datasets, while EEGNN closely matches its performance, demonstrating that early exits incur minimal accuracy trade-off.
[…]
4.3 Empirical Validation of Theoretical Properties
Having evaluated the practical benefits of EEGNN and SAS-GNN in real-world tasks, we now turn to validating the theoretical claims underlying the SAS-GNN block, specifically, its ability to mitigate OST and OSQ according to the theoretical properties expressed in Section 3.
Lower values of $E^{dir}$ indicate smoothing; values near zero signal OST (Cai and Wang, 2020). In Figure 2(a), we report $E^{dir}(\mathbf{H}^{t})$ at each layer on Minesweeper, a heterophilic dataset. GCN rapidly collapses, while GRAFF oversharpens, as described in Di Giovanni et al. (2023). SAS-GNN and A-DGN, in contrast, maintain stable energy profiles, suggesting SAS-GNN inherits its stabilizing behavior primarily from A-DGN.
These trends affirm that SAS-GNN is effective as a proxy to contrast OST and OSQ. Further experimental details, Dirichlet trends, and sensitivity curves appear in Appendix F.2. Appendix F.1 also shows that SAS-GNN stability and non-dissipativeness allow retaining performance even with 100 layers and higher values of $m^{\prime}$ (see Table 8 in the appendix).
(a) $E^{dir}$ on Minesweeper.
[…]
@@ 5 Conclusions
We presented EEGNN, a framework that removes the fixed-depth constraint in GNNs through a differentiable early-exit mechanism at both node and graph levels. At its core lies SAS-GNN, a stable, non-dissipative message-passing scheme designed to mitigate over-smoothing and over-squashing. Empirical results across heterophilic and long-range benchmarks show that EEGNN delivers competitive performance while remaining time and memory-efficient.
Limitations. Despite these advantages, our models may underperform on tasks that demand higher expressive power, likely due to their architectural simplicity. Moreover, we do not directly address the expressiveness limitations of the 1-WL test, nor provide a principled solution to under-reaching.
Future Work. We plan to explore more expressive architectures following the example of this manuscript, by improving the early-exit strategies, and extending EEGNN to edge-level tasks such as link prediction, where exit decisions are more complex due to the pairwise nature of edge representations.
@@ Appendix A Appendix Overview
    We provide information related to GRAFF and A-DGN, which are the foundation of SAS-GNN, but are not discussed in the main paper. We also provide some preliminaries on the Gumbel-softmax distribution.
@@
  • Appendix D: Neural Adaptive-Step Early-Exit for Graph Classification.
    We report the extension of Algorithm 1 for node classification to the task of classifying graphs.
@@
    This section presents extended results from our experiments with EEGNN. We bring additional evidence that SAS-GNN’s design is a proxy to mitigate over-smoothing and over-squashing, and also to retain performance as depth increases. We provide additional baselines on the heterophilic benchmark and LRGB datasets. We provide results on homophilic node classification and graph classification datasets. We test how classic GNNs perform when equipped with our early-exit module. We provide evidence that Early-Exit is a promising direction to significantly improve the GNN performance. We illustrate the nodes’ exit distributions in the discrete and continuous cases that we obtain for the other datasets.
@@ Appendix B Additional Preliminaries
where $\mathbf{\Omega},\mathbf{W}\in\mathbb{R}^{d\times d}$ are trainable matrices, $\sigma$ is a non-linearity, $\mathbf{h}_{i}^{t}$ is the feature of node $i$ at time $t$, $\Gamma(i)$ is the set of neighbors of $i$, $\mathbf{H}^{t}\in\mathbb{R}^{n\times d}$ is the matrix containing all the $d$ dimensional features for each node, $b\in\mathbb{R}^{d}$. The following definitions are from Gravina et al. [2023].
@@
In Gravina et al. [2023], the authors propose an instance of Equation (6) that is both stable and non-dissipative, leading to an evolution of node features that retain all the information collected during the forward pass (e.g., from 00 to an arbitrary $t$). Such an instantiation is
@@
which can be discretized through the Euler method as
@@
As long as $\sigma$ is a non-linearity s.t. $x\sigma(x)\geq 0$ (e.g., ReLU()($\cdot$)( ⋅ ) or tanh()($\cdot$)( ⋅ )), and the weight matrices are symmetric, this rule is proved in Di Giovanni et al. [2023] to minimize the underlying functional:
@@
we consider $a_{i,j}=\frac{1}{\sqrt{d_{i}\cdot d_{j}}}$, which assigns to each edge a weight that depends on the degrees $d_{i}$, $d_{j}$ of the nodes $i$ and $j$. To arrive at this form, we leverage the symmetry of $\mathbf{W}_{s}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$, which allows spectral decomposition as $\mathbf{W}_{s}=\Psi\mathrm{diag}(\boldsymbol{\mu})\Psi^{\top}$. The eigenvalue vector $\boldsymbol{\mu}$ can be split into its positive and negative components, yielding the decomposition:
@@
where $\Theta_{+},\Theta_{-}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$ are lower triangular matrices. Substituting these into the energy functional results in Equation (10), which clearly separates the smoothing (attraction) and sharpening (repulsion) effects. The positive semi-definite part $\Theta_{+}$ contributes to smoothing by encouraging alignment between neighboring node features, while $\Theta_{-}$ induces repulsion, preserving sharp differences. We can see that a gradient flow for such energy, namely Equation (9), can minimize or maximize the edge gradients computed on the node features. This results in a model’s behavior that allows for attraction and repulsion.

### v2

@@ 4.1 Impact of the Early-Exit Components
Model Amazon Ratings $\uparrow$ Tolokers $\uparrow$ Questions $\uparrow$ GCN 51.37 ± 0.34 83.64 ± 0.67 76.58 ± 0.40 SAGE 51.12 ± 0.66 82.43 ± 0.44 76.36 ± 1.50 GAT 51.48 ± 0.28 83.78 ± 0.43 77.95 ± 0.51 GT 51.17 ± 0.66 83.23 ± 0.64 77.95 ± 0.68 Polynormer 54.81 ± 0.49 85.91 ± 0.74 78.92 ± 0.89 Co-GNN 54.17 ± 0.37 84.45 ± 1.17 76.54 ± 0.95 SAS-GNN 51.47 ± 0.68 85.80 ± 0.79 79.60 ± 1.15 EEGNN 51.54 ± 0.5 85.26 ± 0.65 78.90 ± 1.15
Table 4: Performance comparison on ECHO benchmark.
Model diam $\downarrow$ ecc $\downarrow$ sssp $\downarrow$
DRew $1.243\pm 0.047$ $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{4.651\pm 0.020}$} $1.279\pm 0.011$
GraphCON $2.969\pm 0.189$ $5.474\pm 0.001$ $5.734\pm 0.011$
GCN $3.832\pm 0.262$ $5.233\pm 0.034$ $2.102\pm 0.094$
GCNII $2.005\pm 0.093$ $5.241\pm 0.030$ $2.128\pm 0.429$
GIN $1.630\pm 0.161$ $4.869\pm 0.092$ $2.234\pm 0.271$
GPS $2.160\pm 0.098$ $4.758\pm 0.021$ $0.472\pm 0.050$
GRIT $\mathbf{1.014\pm 0.046}$ $5.091\pm 0.158$ $0.121\pm 0.013$
PH-DGN $1.627\pm 0.398$ $5.068\pm 0.126$ $1.323\pm 0.485$
SWAN $1.121\pm 0.070$ $4.840\pm 0.045$ $0.896\pm 0.232$
A-DGN $1.151\pm 0.038$ $4.981\pm 0.037$ $1.176\pm 0.140$
GRAFF $2.731\pm 0.052$ $5.344\pm 0.086$ $3.363\pm 1.356$
APGCN - $11.080\pm 0.099$ $6.729\pm 0.018$
SAS-GNN $1.914\pm 0.397$ $4.806\pm 0.018$ $0.244\pm 0.243$
EEGNN $1.805\pm 0.068$ $5.048\pm 0.237$ $\mathbf{0.065\pm 0.074}$
4.2 Comparison with the Heterophilic and ECHO Benchmarks
Datasets. (i) Heterophilic Node Classification. We use Amazon Ratings (ACC), Tolokers (AUROC), and Questions (AUROC(Platonov et al., 2023) to assess backbone robustness in low-homophily settings, where connected nodes frequently differ in labels. Our goal is to verify if our parameter-efficient backbone can match state-of-the-art methods in these structurally complex scenarios.
(ii) The ECHO Benchmark. To evaluate long-range capabilities, we employ the recently released ECHO benchmark (Miglior et al., 2025). This suite was designed to bridge the gap left by previous benchmarks, such as LRGB (Dwivedi et al., 2023), where tasks were often solvable with shallow receptive fields. We report performance on the synthetic node-level tasks sssp and ecc, and the synthetic graph-level task diam, which require several hop of message-passing in order to perform well. This benchmark provides an ideal setting to validate the adaptivity of our early-exit mechanism. Specifically, we aim to test how EEGNN ranks among long-range specialized architectures, and how it 1) correctly identifies the depth required by the task node/graph-wise, and 2) keeps its features reliable for inference, without suffering from depth-related degradation. Furthermore, we highlight that the ECHO benchmark encompasses a wide and heterogeneous set of topologies (Miglior et al., 2025); these structures frequently induce topological bottlenecks, which typically cause performance degradation via over-squashing (Arnaiz-Rodriguez and Errica, 2025), presenting another crucial challenge for EEGNN. In the appendix, we provide additional results on the original LRGB datasets, transductive homophilic node classification, graph classification (TUDataset (Morris et al., 2020)), and large-scale OGB benchmarks (Hu et al., 2021) in Appendix F.9F.11. Full dataset descriptions and implementation details are provided in Appendix E.
Baselines. For heterophilic datasets (Platonov et al., 2023), we compare with standard MPNNs (GCN, SAGE, GAT), asynchronous methods (Co-GNN), and Transformers (GT, Polynormer). Crucially, to ensure a fair comparison with our theoretically constrained backbone, we adopt MPNN variants without dropout or normalization, as they are also the main responsible for boosting performance (Luo et al., 2024). For the ECHO benchmark, we categorize baselines into four groups: (i) standard MPNNs (GCN, GIN (Xu et al., 2019), GCNII (Chen et al., 2020)); (ii) rewiring methods as DRew (Gutteridge et al., 2023); (iii) Graph Transformers (GPS (Rampášek et al., 2023), GRIT (Ma et al., 2023)); and (iv) Graph NODEs (GraphCON (Rusch et al., 2022), A-DGN, SWAN (Gravina et al., 2025), PH-DGN (Heilig et al., 2024)), the same class of SAS-GNN and EEGNN. Additionally, we include APGCN (Spinelli et al., 2021), to represent budget driven early-exit GNNs. We also implement GRAFF (Di Giovanni et al., 2023) as a form of symmetric-only ablation, complementing the antisymmetric dynamics of A-DGN. Full details on hyperparameters are provided in Appendix E.4.
Results. On the heterophilic datasets (Table 3), both SAS-GNN and EEGNN consistently match or surpass standard MPNN baselines and frequently outperform more complex GTs. For instance, in Questions, SAS-GNN achieves state-of-the-art performance, while EEGNN maintains robust accuracy across all tasks. It is important to note that for GCN, SAGE, and GAT, we report results without dropout. This ensures a fair comparison, as both SAS-GNN and EEGNN are trained without normalization layers, dropout, or bias terms. This design choice is imposed to satisfy the theoretical conditions of our stability theorems, and minimization of the energy functional. We further analyze the efficiency advantages resulting from this lightweight design later in this section.
The most significant capabilities of our proposed framework are evident in the ECHO benchmark (Table 4).Here, EEGNN and SAS-GNN track closely with GraphNODEs explicitly specialized for long-range reasoning (e.g., SWAN, A-DGN, PH-DGN). Notably, in the sssp task, both models rank among the top-3, with EEGNN achieving the best overall performance (MAE 0.065), significantly outperforming the runner-up GRIT (0.121). In other metrics, our models generally offer a superior alternative to standard MPNNs (with the minor exception of GIN on diam).
Unlike standard MPNNs (e.g., GCN), EEGNN manages to outperform GTs on specific tasks. This result is particularly significant given that GTs are inherently resilient to topological over-squashing due to their global attention mechanisms. The superior performance of EEGNN suggests that the inductive biases of the SAS-GNN backbone effectively mitigate topological bottlenecks, allowing for stable deep propagation, while the exit module successfully identifies a meaningful halting point, as in Figure 2. Finally, the comparison with APGCN is particularly instructive. As shown in Table 4, APGCN exhibits always high test MAE. This failure stems primarily from its budget-aware halting mechanism, which forces nodes to exit prematurely. In particular, when long-range reasoning is required, harder samples are ignored rather than deeply processed. In contrast, EEGNN’s success corroborates the importance of our Neural Adaptive Step: by learning exit decisions directly via task loss gradients, our method ensures that the chosen halting points align strictly with the computational needs of the task.
4.3 Accuracy–Efficiency Trade-off
While our primary objective is not to surpass the raw accuracy of large models, EEGNN aims to occupy a distinct operational point: delivering competitive performance with superior parameter efficiency. To assess this, we analyze the trade-off between accuracy and model size (trainable parameters) using the best-performing hyperparameters from Table 3. Figure 5 illustrates this relationship on the Questions dataset. Thanks to the layer-wise weight sharing inherited from the SAS-GNN backbone, EEGNN maintains a constant parameter count regardless of depth. This allows it to reach strong accuracy with an order of magnitude fewer parameters than heavier architectures such as Polynormer, SGFormer (Wu et al., 2024) or NodeFormer (Wu et al., 2023). Although EEGNN may slightly underperform on datasets where dropout or normalization are essential for generalization, it may offer a highly favorable balance for resource-constrained scenarios. By avoiding the parameter explosion typical of deep GNNs, EEGNN proves that encompassing adaptive exits can be integrated without inflating the computational budget. Additional analyses, including inference time comparisons, are provided in Appendix F.7.
Figure 5: Accuracy-Efficiency trade-off on Questions.
4.4 Practical Recommendations
When to use. EEGNN is ideal when the required reasoning depth is unknown or variable across samples (e.g., ECHO tasks). Its success relies on a core design principle: early exits are only reliable when the backbone resists deep processing. Without a stable backbone like SAS-GNN, premature halting often acts merely as a safeguard against OST or topological OSQ rather than a confident decision. EEGNN combines stability with adaptivity, ensuring that the model continues processing until it is truly confident, but without requiring full-depth processing.
When not to use. As a Graph NODE, EEGNN inherits certain limitations. It tends to lag behind heavily regularized architectures or with a large number of parameters on tasks where such mechanisms are critical for generalization. In these cases, where the strict theoretical constraints of ODEs (no dropout/bias) limit performance, or if the problem radius is self-contained, we recommend using other models.
Hyperparameter strategy. Unlike fixed-depth GNNs where depth ($L$) is a critical hyperparameter requiring extensive tuning, EEGNN treats $L$ as a budget. Since our results (Figure 3) show that performance does not degrade with increased depth, we recommend setting $L$ to the maximum value allowed by hardware constraints. This simplifies the search space, allowing practitioners to focus tuning efforts on other hyperparameters, trusting the Neural Adaptive Step to identify the effective depth per sample.
@@ 5 Conclusions
We presented EEGNN, an end-to-end framework that removes the fixed-depth limitation of conventional GNNs by introducing a differentiable early-exit mechanism at both node and graph levels. This design enables the network to adapt its depth dynamically based on input complexity: simple inputs exit early to conserve computation, while complex inputs, requiring long-range propagation or richer feature extraction, traverse deeper layers. This flexibility allows users to tailor the model fixing the number of layers as a budget, optimizing for either strict resource budgets or maximum reasoning capacity. Unlike previous early-exit approaches, which often ignore the inherent instability of deep MPNNs, EEGNN explicitly addresses this challenge through SAS-GNN. This novel, provably stable backbone ensures that intermediate representations remain informative, allowing the model to safely leverage deep propagation when necessary without suffering from classic depth-related message-passing flaws. Empirical evaluation on heterophilic and long-range benchmarks confirms that EEGNN achieves competitive accuracy, while maintaining constant memory complexity and superior runtime efficiency, while adapting the exit choices to the task at hand.
Limitations. Despite these advantages, EEGNN inherits some limitations of Graph Neural ODE–style architectures, potentially underperforming on tasks requiring extremely high expressive power. Furthermore, EEGNN does not inherently overcome the expressiveness limits of the 1-WL test, nor does it offer a formal solution to under-reaching issues.
Future Work. Future directions include: (i) integrating more expressive backbones that retain stability in deep configurations, and (ii) extending the early-exit mechanism to edge-level tasks (e.g., link prediction), where handling pairwise dependencies presents a non-trivial challenge.
@@ Appendix A Appendix Overview
    We provide information related to GraphNODEs, GRAFF and A-DGN, which are the foundation of SAS-GNN, but are not discussed in the main paper. We discuss the seminal work on early-exit mechanisms for GNNs, which is APGCN, that we also use in this paper. We also provide some preliminaries on the Gumbel-softmax distribution.
@@
  • Appendix D: Additional details on Neural Adaptive-Step Early-Exit.
    We report the extension of Algorithm 1 for node classification to the task of classifying graphs. We comment on how EEGNN is related to the underreaching phenomenon, and we draw a connection of Neural Adaptive Step with Adaptive-Step Runge-Kutta solvers.
@@
    This section presents extended results from our experiments with EEGNN. We bring additional evidence that SAS-GNN’s design is a proxy to mitigate over-smoothing and over-squashing, and also to retain performance as depth increases. We provide additional baselines on the heterophilic benchmark and LRGB datasets. We provide results on short-range, small, and large-scale homophilic node classification and graph classification datasets. We test how classic GNNs perform when equipped with our early-exit module. We provide evidence that Early-Exit is a promising direction to significantly improve the GNN performance. We illustrate the nodes’ exit distributions in the discrete and continuous cases that we obtain for the other datasets.
@@ Appendix B Additional Preliminaries
In this section, we provide the necessary background on the specific architectures that inspire our framework, namely Graph Neural ODEs (GraphNODEs), A-DGN, and GRAFF. Furthermore, we detail the Adaptive Propagation GCN (APGCN), the seminal work on early exiting for graphs, and the Gumbel-Softmax distribution, which enables the differentiability of our discrete exit decisions.
Graph Neural Ordinary Differential Equations (GraphNODEs). GraphNODEs [Poli et al., 2021] generalize discrete Graph Neural Networks (GNNs) by modeling the evolution of node features $\mathbf{H}(t)$ as a continuous-time dynamical system. Instead of a fixed sequence of layers $l=1,\dots,L$, the feature transformation is governed by an Ordinary Differential Equation (ODE):
$\frac{d\mathbf{H}(t)}{dt}=f(\mathbf{H}(t),\mathcal{G},\theta_{t}),\quad\mathbf{H}(0)=\mathbf{X},$ (5)
where $\mathbf{X}$ represents the input features, $\mathcal{G}$ the graph structure, and $f$ is a neural network (e.g., a GNN layer) parametrized by $\theta$. The output features are obtained by integrating this system up to a time $T$:
$\mathbf{H}(T)=\mathbf{H}(0)+\int_{0}^{T}f(\mathbf{H}(t),\mathcal{G},\theta_{t})dt.$ (6)
This framework allows the use of many ODE solvers (e.g., Euler, Runge-Kutta) for the forward pass and adjoint methods for memory-efficient backpropagation. If we consider the Euler integration (i.e. what we did in the main text) with step size $\tau$, Equation (6) becomes the following iterative update for each step $k$:
$\mathbf{H}^{k+1}=\mathbf{H}^{k}+\tau f(\mathbf{H}^{k},\mathcal{G},\theta_{k}).$ (7)
Our proposed EEGNN follows this continuous-depth paradigm but introduces a mechanism to dynamically determine the integration endpoint $T_{i}$ for each node or graph, which relies on considering $\tau$ not as fixed, but as the output of a neural network. In the main text we refer to it as Neural Adaptive Step.
@@
where $\mathbf{\Omega},\mathbf{W}\in\mathbb{R}^{d\times d}$ are trainable matrices, $\sigma$ is a non-linearity, $\mathbf{h}_{i}^{t}$ is the feature of node $i$ at time $t$, $\Gamma(i)$ is the set of neighbors of $i$, $\mathbf{H}^{t}\in\mathbb{R}^{n\times d}$ is the matrix containing all the $d$ dimensional features for each node, $b\in\mathbb{R}^{d}$. The following definitions are from [Gravina et al., 2023].
@@
In [Gravina et al., 2023], the authors propose an instance of Equation (8) that is both stable and non-dissipative, leading to an evolution of node features that retain all the information collected during the forward pass (e.g., from 0 to an arbitrary $t$). Such an instantiation is
@@
which can be discretized through the Euler method (i.e. This is because it follows the GraphNODE’s paradigm) as
@@
As long as $\sigma$ is a non-linearity s.t. $x\sigma(x)\geq 0$ (e.g., ReLU()($\cdot$) or tanh()($\cdot$)), and the weight matrices are symmetric, this rule is proved in [Di Giovanni et al., 2023] to minimize the underlying functional:
@@
We consider $a_{i,j}=\frac{1}{\sqrt{d_{i}\cdot d_{j}}}$, which assigns to each edge a weight that depends on the degrees $d_{i}$, $d_{j}$ of the nodes $i$ and $j$. We leverage the symmetry of $\mathbf{W}_{s}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$, which allows spectral decomposition as $\mathbf{W}_{s}=\Psi\mathrm{diag}(\boldsymbol{\mu})\Psi^{\top}$. The eigenvalue vector $\boldsymbol{\mu}$ can be split into its positive and negative components, yielding the decomposition:
@@
where $\Theta_{+},\Theta_{-}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$ are lower triangular matrices.
We can see that a gradient flow for such energy, namely Equation (11), can minimize or maximize the edge gradients computed on the node features. This results in a model’s behavior that allows for attraction and repulsion weighted by $\Theta_{+}$, and $\Theta_{-}$, specifically the positive semi-definite part $\Theta_{+}$ contributes to smoothing by encouraging alignment between neighboring node features, while $\Theta_{-}$ induces repulsion, preserving sharp differences.
Adaptive Propagation GCN (APGCN). APGCN [Spinelli et al., 2021] is the first approach to introduce adaptive depth in GNNs using a halting mechanism inspired by Adaptive Computation Time (ACT) [Graves, 2017] for RNNs. In APGCN, each node $i$ maintains a halting probability $p_{i}^{k}$ at each layer $k$. The halting unit is defined as a sigmoid function over the current node state: $h_{i}^{k}=\sigma(\text{MLP}(\mathbf{h}_{i}^{k}))$. The process continues until the cumulative probability exceeds a threshold:
$N(i)=\min\{k^{\prime}:\sum_{k=1}^{k^{\prime}}h_{i}^{k}\geq 1-\epsilon\},$ (15)
where $1-\epsilon$ is a confidence threshold (typically 0.99). Crucially, APGCN requires an auxiliary ponder cost added to the loss function, $\mathcal{L}_{budget}=\sum_{i}N(i)+R_{i}$, to penalize the number of steps and prevent the model from running indefinitely. This explicit penalty creates a "budget-aware" bias, where the model is incentivized to exit early to minimize the loss, potentially at the cost of task performance (as discussed in Section 4.1). The final loss for APGCN becomes:
$\mathcal{L}=\mathcal{L}_{task}+\mu\cdot\mathcal{L}_{budget}$ (16)
Here $\mu$ is a coefficent needed to weight the need for budget in the early-exit setting. We remark that the halting unit cannot be trained end-to-end if this additional loss is missing.